At DMV, there are 4 service windows handling new license applications. Each window has one employee. On average, it takes one employee 6 minutes to complete one application.

The demand rate of the new license application is 25 applications/hour.

Please use the information provided to answer the next 3 questions.

6.) Accounting all 4 service windows, the stage capacity of the new license application service is ________ applications/hour.

7.) The flow rate of the new license application service is ___________ applications/hour.

8.) The cycle time of the license application service is ________ minutes/application?
Answer #1

6

Stage capacity = Number of resources / Processing time per unit = (4/6) applications per minute = 4*60/6 applications per hour = 40 applications per hour.

7.

Demand rate = 25 applications per hour.

Flow rate = Min{Capacity, Demand} = Min(40, 25) = 25 application per hour.

8.

Cycle time = 1 / Flow rate = (1/25) hour = 2.5 minutes